OpenStudy (anonymous):

Sharon has 8 cups of granola. She eats cup each day. How long will her granola last?

OpenStudy (anonymous):

She eats 2/3 of a cup.

OpenStudy (danjs):

8 - (number of days eating) x (cups eaten per day) = 0 left

OpenStudy (danjs):

so 8 equals, number of days eating times 2/3 cups each day

OpenStudy (anonymous):

Thanks

OpenStudy (danjs):

you can divide 8 by 2/3 , or set up a little thing like this 8 = (2/3)*days

OpenStudy (danjs):

3*8 = 2days 24 = 2 days 24/2 = days 12 = days

OpenStudy (anonymous):

That makes so much more sense!!

OpenStudy (danjs):

are you in algebra, or before algebra?

OpenStudy (anonymous):

In algebra, I'm studying for a test, I'm homeschooled.

OpenStudy (danjs):

oh , ok, let d=days eating total = #days x cups per day 8 = (2/3)*d

OpenStudy (danjs):

solve for d

OpenStudy (danjs):

that is what i did above. d=12 days